The locomotive passes through the station at a speed of 16 m / s. When leaving the station, the locomotive accelerates with a constant acceleration of 0.33 m / s until it reaches a speed of 35 m / s. Determine the path that the locomotive will travel during acceleration.


1
Expert's answer
2021-03-21T11:24:25-0400

We can find the path that the locomotive will travel during acceleration as follows:


"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a},""d=\dfrac{(35\ \dfrac{m}{s})^2-(16\ \dfrac{m}{s})^2}{2\cdot0.33\ \dfrac{m}{s^2}}=1468\ m."